OpenStudy (anonymous):

can some one help me -24x^2+28/3x-5/6=0

OpenStudy (anonymous):

\[\frac{ -24x^2 + 28 }{ 3x - 5 } * \frac{ 1 }{ 6 } = 0\] it look like that?

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

the problem looks like this -24x^2+28/3x-5/6=0

OpenStudy (anonymous):

I took the reciprocal of 6 for it's being divide by another fraction. (if you divide by a fraction you flip it and multiply)

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

multiply \[\frac{ -24x^2 + 28 }{ 3x - 5 } * \frac{ 1 }{ 6 } = 0\] and what does it look like?